OpenStudy (anonymous):

2/3 of something is?

sammixboo (sammixboo):

Depends on what that something

OpenStudy (anonymous):

of what?

OpenStudy (dmndlife24):

2/3 of something is? Let "something be x" \[2/3 \times x=\] 2/3x :)

OpenStudy (anonymous):

Based on the figure, Jordan says that 2/3RW=RU. Kendra disagrees. Is either of them correct? RW=20 imches RU=30 imches

OpenStudy (anonymous):

Jordan would be wrong and Kendra would be right. Because she should have done 2/3RU = RW